# Causal Inference: A Tale of Three Frameworks

### v1

@@ 1 Introduction
Causal inference is the science of understanding the consequences of interventions, requiring assumptions that extend beyond those needed for purely associational analysis. Its importance has grown rapidly in the era of machine learning and artificial intelligence, where the ability to draw reliable causal conclusions is central to building systems that are not only predictive but also trustworthy, transparent, and robust to distributional shifts (Peters et al.,, 2016; Wachter et al.,, 2017; Pearl,, 2019; Arjovsky et al.,, 2020; Bühlmann,, 2020; Tjoa and Guan,, 2021; Schölkopf,, 2022; Jiao et al.,, 2024). Over the past decades, three foundational frameworks have emerged to formalize causal reasoning: the potential outcomes framework, nonparametric structural equation models (NPSEMs), and directed acyclic graphs (DAGs). Each framework carries its own formal machinery, conceptual underpinnings, and historical roots. Although they originated in distinct disciplinary traditions, they are now increasingly recognized as complementary, and in many cases translatable into one another.
@@
A substantial literature surveys causal inference from within a single framework or with an emphasis on identification and estimation approaches (e.g. Imbens and Wooldridge,, 2009; Spirtes,, 2010; Pearl, 2010a, ; Kuang et al.,, 2020; Yao et al.,, 2021; Li et al.,, 2023; Jiao et al.,, 2024). However, there has been little concise, side-by-side treatment that translates assumptions and results across the three frameworks and clarifies when they agree and when they differ. This review aims to fill that gap. We offer a selective review of the three frameworks, with the goal of elucidating their differences, similarities, and interconnections. Rather than offering a comprehensive survey, we aim to guide readers, particularly those new to the field or approaching causal inference from a single perspective, through the conceptual landscape of causal inference using simple, illustrative examples.
@@
In this review, we focus on interventional questions, which ask what would happen under specific actions or treatments applied to the entire population, without conditioning on what actually occurred. These contrast with counterfactual questions, which ask what would have happened for the same individual, had things been different, given what actually occurred. For example, as noted by Peters et al., (2017, Remark 6.20(ii)), suppose someone offers you \$10,000 if you correctly predict the result of a coin flip. You guess “heads” and lose. The interventional effect of guessing “heads” versus “tails” is zero, whereas the counterfactual effect, given that you actually guessed “heads” and lost, is \$10,000. While counterfactuals play a central role in philosophical and legal reasoning, their identification typically requires stronger, often untestable, assumptions. See Remark 3 for further discussion of this distinction.
@@
Although we do not delve into estimation methods or unmeasured confounding in depth, we briefly comment on these in later sections to clarify their connection to the foundational frameworks. Other approaches, such as Dawid’s decision-theoretic framework (Dawid,, 2000, 2015; Richardson and Robins,, 2023), are also omitted from our discussion. The decision-theoretic perspective defines causality through conditional independence relations among decisions, observations, and consequences, without invoking counterfactuals or structural equations. Conceptually, it shares common ground with all three frameworks reviewed here: like potential outcomes, it distinguishes between observed and hypothetical actions; like structural models, it emphasizes invariance of mechanisms under intervention; and like causal DAGs, it encodes causal assumptions through independence relations. However, it differs from these frameworks in that it formulates causal reasoning entirely within the observable world, avoiding reference to unobservable counterfactual quantities. Because our focus is on frameworks that explicitly model potential outcomes or interventions, we do not explore the decision-theoretic approach in detail. For these topics, and for comprehensive treatments of any individual framework, we refer readers to existing textbooks (e.g., Pearl,, 2009; Peters et al.,, 2017; Imbens and Rubin,, 2015; Hernán and Robins,, 2025) and the broader specialized literature.
@@
The frameworks reviewed here primarily concern single-stage interventions. Many applied settings, however, involve sequential or adaptive decisions, such as treatment policies in medicine, dynamic pricing, or reinforcement learning. These settings have been extensively studied in the literature on dynamic treatment regimes and longitudinal causal inference, most notably through Robins’s development of g-methods and related approaches (Robins,, 1986; Hernán and Robins,, 2025). These frameworks generalize the ideas discussed here to multi-stage interventions and time-varying treatments. Because our focus is on the conceptual connections among the three foundational frameworks, we do not cover these extensions in detail.
@@ 2.1 The Potential Outcomes Framework
The potential outcomes framework provides an intuitive and mathematically tractable foundation for causal inference. It was originally introduced by Neyman, (1923) for randomized experiments and was later formalized and extended to observational studies by Rubin, (1974). This framework is closely related to an independent line of work in econometrics that developed the so-called switching regression model (Roy,, 1951; Manski,, 1993). These developments allow researchers to pose well-defined causal questions, even in the absence of randomized experiments, by explicitly invoking assumptions that link observed data to potential outcomes. We refer interested readers to Rubin, (2005); Morgan and Winship, (2014); Imbens and Rubin, (2015); Hernán and Robins, (2025) for a more detailed introduction.
@@
The variables $Y(0)$ and $Y(1)$ are two distinct random variables representing hypothetical outcomes under different treatment assignments. They are not different realizations of the same variable. For each unit, only one of these two outcomes is realized and observed, depending on the actual treatment received. The other remains counterfactual. This leads to the Fundamental Problem of Causal Inference (Holland,, 1986): we can never simultaneously observe both potential outcomes $Y(0)$ and $Y(1)$ for the same individual. From this perspective, the potential outcomes framework views causal inference as a missing data problem, since at least one of the two potential outcomes is missing for each unit.
@@
Potential outcomes are sometimes referred to as counterfactuals in the literature. Before treatment assignment, both $Y(1)$ and $Y(0)$ are potential outcomes. After treatment assignment, the outcome corresponding to the received treatment is factual, while the other is counterfactual. To avoid ambiguity, we use the term potential outcomes to collectively refer to both $Y(1)$ and $Y(0)$.
@@
The estimands we have introduced so far are sometimes referred to as interventional estimands, as they pertain to the distribution of outcomes under hypothetical interventions, such as setting the treatment to a fixed value. Another type of question that may be of interest involves counterfactual reasoning. For example, when $Y$ is binary, the persuasion rate is defined as $\mathbb{P}\{Y(1)=1\mid Y(0)=0\}$ (Jun and Lee,, 2023), which quantifies the probability that an individual would take an action under treatment given that they would not have done so under control. Other examples include the probability of necessity, $\mathbb{P}\{Y(0)=0\mid A=1,Y=1\}$; the probability of sufficiency, $\mathbb{P}\{Y(1)=1\mid A=0,Y=0\}$; and the probability of necessity and sufficiency, $\mathbb{P}\{Y(1)=1,Y(0)=0\}$ (Pearl,, 2009, § 9.2). Identification of counterfactual estimands is typically much more involved, as it relies on the dependence between potential outcomes, which cannot be recovered from observed data, even in randomized controlled experiments. We refer readers to Peters et al., (2017, Example 3.4) for an interesting example and discussions on whether counterfactuals or interventional quantities should serve as the basis for legal decisions; see also Pearl, (2009, §1.4.4). Along with association questions, interventional and counterfactual questions are referred to as the three levels of the ladder of causation by Pearl and Mackenzie, (2018, § 1).
A crucial component of the Potential Outcome framework is the Stable Unit Treatment Value Assumption (SUTVA) (Rubin,, 1980), which posits that (i) there are no multiple versions of the treatment, and (ii) one unit’s outcome is unaffected by the treatment assigned to another unit, meaning that there is no interference between units.
The first part of SUTVA is closely related to the consistency assumption (Cole and Frangakis,, 2009; VanderWeele,, 2009; Pearl, 2010b, ), which is the fundamental assumption that links the observational world to the potential outcomes in a hypothetical interventional world. The consistency assumption states that if an individual receives treatment level $A=a$, then their observed outcome $Y$ equals their potential outcome under that treatment:
@@
A classic example of an ill-defined treatment arises when the so-called “treatment” is the causal effect of obesity (defined as BMI>30>30) on heart disease (Hernán and Robins,, 2025).
@@
The consistency assumption discussed here is not the same as the notion of consistency used to describe a statistical property of an estimator (e.g., Lehmann and Casella,, 1998), which refers to convergence in probability to the true parameter value.
The second part of SUTVA, namely the no interference assumption, implies that each unit’s potential outcome depends only on its own treatment assignment and not on the treatment received by others. This assumption can be violated in many real-world contexts. For example, in vaccine studies, the likelihood of an individual contracting an infectious disease may depend not only on whether they are vaccinated but also on the vaccination status of others in their community. Such interference is common in infectious disease epidemiology and social network settings, where spillover, peer effects, or herd immunity can induce dependencies across units. A growing body of work has extended the potential outcomes framework to accommodate interference. We refer interested readers, for example, to Kuang et al., (2020, §9) and Bajari et al., (2023) for surveys of this area.
@@
This condition implies that, conditional on $L$, the treatment assignment $A$ is as good as random and thus independent of the potential outcomes. Under this assumption and a positivity condition that ensures all levels of treatment are possible across values of $L$, the ACE is identified by the following g-formula (Robins,, 1986):
@@
where the outer expectation is over the marginal distribution of $L$. This expression can be estimated using regression, stratification, matching, inverse probability weighting, or doubly robust methods, depending on the structure and quality of the observed data.
@@ 2.2 Structural Equation Models
An NSPEM (Haavelmo,, 1943; Strotz and Wold,, 1960; Pearl,, 2009; Spirtes et al.,, 2000; Halpern and Pearl,, 2000) consists of a finite set of random variables $V=\{V_{1},\ldots,V_{p}\}$, a collection of exogenous (unobserved) random variables $\varepsilon=(\varepsilon_{1},\ldots,\varepsilon_{p})$, and a collection of measurable functions $f_{1},\ldots,f_{p}$, such that each endogenous variable $V_{j}$ is determined by an equation of the form
@@
Together, the functions $f_{j}$ and exogenous variables $\varepsilon_{j}$ specify a data-generating process and thereby induce a joint distribution over $V$. Importantly, NPSEMs do not need to be acyclic: cycles or feedback relations can in principle be allowed. When the system is acyclic, however, there exists (at least) one ordering $V_{\pi(1)},\ldots,V_{\pi(p)}$ such that each $V_{\pi(j)}$ depends only on variables earlier in the ordering. In this case, the structure of dependence among variables can be represented by a directed acyclic graph (DAG) (see Section 3.2). The model is then fully defined by the tuple $(f_{1},\ldots,f_{p},\varepsilon_{1},\ldots,\varepsilon_{p})$ together with a valid causal ordering (e.g., as obtained from Simon’s causal ordering algorithm (Simon,, 1953)). The nonparametric nature of NPSEMs arises from the fact that no parametric form is assumed for the functions $f_{j}$, allowing for considerable flexibility in modeling complex causal relationships. A key assumption in many NPSEM applications is that the exogenous variables in $\varepsilon$ are jointly independent, yielding the special case known as the NPSEM with Independent Errors (NPSEM-IE).
@@
The Causally Interpretable Structured Tree Graph (CISTG) as detailed as the data (Robins,, 1986) is mathematically equivalent to an NPSEM model (with no assumption on the errors); see Equations (10) and (14) below.
@@
The key assumption in the SEM framework that links the observational world to interventional worlds is the autonomy assumption (Aldrich,, 1989), also known as the invariance or modularity assumption (Peters et al.,, 2017). It asserts that each structural equation represents an independent, self-contained causal mechanism that remains invariant under interventions on other variables. Formally, an intervention on a variable $A$ corresponds to removing the structural equation for $A$ and replacing it with a constant assignment. This is denoted by Pearl’s do-operator (Pearl,, 2009), written as $\text{do}(A=a)$. Historically, Robins, (1986) used the notation $g=a$ to express the same idea; Pearl’s do notation is now more widely used.
@@
Autonomy ensures that interventions can be represented by replacing the structural equation for the intervened variable, without having to re-specify the rest of the system. This feature allows us to predict counterfactual outcomes under hypothetical interventions. In real-world systems, however, autonomy is not always guaranteed (e.g., Cartwright,, 2007). Administering the treatment may change the patient’s behavior or alter how the outcome is measured. For instance, it may modify the reporting mechanism or induce side effects that feed back into the outcome-generation process in an unmodeled way. In such cases, the function $f_{Y}$ may no longer be invariant to interventions on $A$. In other words, the mapping from $(L,A,\varepsilon_{Y})$ to $Y$ may differ depending on whether $A$ is set naturally or via intervention.
@@
The term Structural Causal Models (SCMs) is often used in the literature. It may refer to the NPSEM (e.g., Pearl,, 2009, p. 203, Definition 7.1.1) alone or specifically to the NPSEM with independent errors (NPSEM-IE) (e.g., Pearl,, 2009, p. 44, Definition 2.2.2); see also Peters et al., (2017, p. 62, Definition 6.2). Following Shpitser et al., (2022), we use our terminology here to emphasize the distinction between NPSEM and NPSEM-IE.
@@
The independent error and autonomy assumptions in the NPSEM-IE framework imply the principle of independent mechanisms (PIM) (Janzing and Schlölkopf,, 2010; Peters et al.,, 2017). This principle states that the causal generative process can be decomposed into independent modules, one for each variable, that remain invariant to changes in the distributions of other variables. In the special case of two variables, PIM reduces to the so-called independence of cause and mechanism, which asserts that the conditional distribution $P(Y\mid X)$ is independent of the marginal distribution $P(X)$. In other words, under NPSEM-IE assumptions, the causal mechanism mapping $X\mapsto Y$ is invariant to changes in the distribution of $X$.

### v2

@@ 1 Introduction
Causal inference is the science of understanding the consequences of interventions, requiring assumptions that extend beyond those needed for purely associational analysis. Its importance has grown rapidly in the era of machine learning and artificial intelligence, where the ability to draw reliable causal conclusions is central to building systems that are not only predictive but also trustworthy, transparent, and robust to distributional shifts (Peters et al., 2016; Wachter et al., 2017; Pearl, 2019; Arjovsky et al., 2020; Bühlmann, 2020; Tjoa and Guan, 2021; Schölkopf, 2022; Jiao et al., 2024). Over the past decades, three foundational frameworks have emerged to formalize causal reasoning: the potential outcomes framework, nonparametric structural equation models (NPSEMs), and directed acyclic graphs (DAGs). Each framework carries its own formal machinery, conceptual underpinnings, and historical roots. Although they originated in distinct disciplinary traditions, they are now increasingly recognized as complementary, and in many cases translatable into one another.
@@
A substantial literature surveys causal inference from within a single framework or with an emphasis on identification and estimation approaches (e.g. Imbens and Wooldridge, 2009; Spirtes, 2010; Pearl, 2010a; Kuang et al., 2020; Yao et al., 2021; Li et al., 2023; Jiao et al., 2024). However, there has been little concise, side-by-side treatment that translates assumptions and results across the three frameworks and clarifies when they agree and when they differ, despite earlier comparative discussions such as Greenland and Brumback (2002), Pearl (2009), and Robins and Richardson (2011). This review aims to fill that gap. We offer a selective review of the three frameworks, with the goal of elucidating their differences, similarities, and interconnections. Rather than offering a comprehensive survey, we aim to guide readers, particularly those new to the field or approaching causal inference from a single perspective, through the conceptual landscape of causal inference using simple, illustrative examples.
@@
In this review, we focus on interventional questions, which ask what would happen under specific actions or treatments applied to the entire population, without conditioning on what actually occurred. These contrast with counterfactual questions, which ask what would have happened for the same individual, had things been different, given what actually occurred. For example, as noted by Peters et al. (2017, Remark 6.20(ii)), suppose someone offers you \$10,000 if you correctly predict the result of a coin flip. You guess “heads” and lose. The interventional effect of guessing “heads” versus “tails” is zero, whereas the counterfactual effect, given that you actually guessed “heads” and lost, is \$10,000. While counterfactuals play a central role in philosophical and legal reasoning, their identification typically requires stronger, often untestable, assumptions. See Remark 3 and Wu and Wang (2026) for further discussion of this distinction.
@@
Although we do not delve into estimation methods or unmeasured confounding in depth, we briefly comment on these in later sections to clarify their connection to the foundational frameworks. Other approaches, such as Dawid’s decision-theoretic framework (Dawid, 2000, 2015; Richardson and Robins, 2023), are also omitted from our discussion. The decision-theoretic perspective defines causality through conditional independence relations among decisions, observations, and consequences, without invoking counterfactuals or structural equations. Conceptually, it shares common ground with all three frameworks reviewed here: like potential outcomes, it distinguishes between observed and hypothetical actions; like structural models, it emphasizes invariance of mechanisms under intervention; and like causal DAGs, it encodes causal assumptions through independence relations. However, it differs from these frameworks in that it formulates causal reasoning entirely within the observable world, avoiding reference to unobservable counterfactual quantities. Because our focus is on frameworks that explicitly model potential outcomes or interventions, we do not explore the decision-theoretic approach in detail. For these topics, and for comprehensive treatments of any individual framework, we refer readers to existing textbooks (e.g., Pearl, 2009; Peters et al., 2017; Imbens and Rubin, 2015; Hernán and Robins, 2025) and the broader specialized literature.
Another influential framework that we do not review in detail is the sufficient–component cause (SCC) framework; see Rothman (1976), Greenland and Poole (1988), Greenland and Brumback (2002), Rothman et al. (2008, Chap. 2), VanderWeele and Robins (2009), VanderWeele and Richardson (2012, §7), and VanderWeele (2015, p.10 & Chap. 10) for detailed discussions. SCC models provide a mechanistic representation of causation and are well suited for questions about explanatory causal structures and mechanistic interaction. These mechanistic questions differ somewhat from the interventional questions that motivate our review, and several SCC quantities cannot be expressed within the causal DAG formalism. Since our goal here is to compare three frameworks for interventional causal questions, we do not cover SCC models, although they remain an important component of the broader causal inference literature.
A further set of traditions, especially influential in health and legal applications, involves more informal approaches to causal inference alongside formal models. Greenland (2004), for example, distinguishes canonical approaches based on considerations such as the Hill’s list (Hill, 1965), in which causality is treated as a property of an observed association to be judged using qualitative “symptoms” like temporality, strength, consistency, and dose response, from methodologic modeling or bias analysis, which uses parametric models and sensitivity analyses to assess the impact of selection, unmeasured confounding, and measurement error on effect estimates. These approaches are widespread in applied practice and can be combined with potential outcome, structural, and graphical models, but they fall outside our focus on three foundational frameworks for representing and identifying interventional effects.
@@
The frameworks reviewed here primarily concern single-stage interventions. Many applied settings, however, involve sequential or adaptive decisions, such as treatment policies in medicine, dynamic pricing, or reinforcement learning. These settings have been extensively studied in the literature on dynamic treatment regimes and longitudinal causal inference, most notably through Robins’s development of g-methods and related approaches (Robins, 1986; Hernán and Robins, 2025). These frameworks generalize the ideas discussed here to multi-stage interventions and time-varying treatments. Because our focus is on the conceptual connections among the three foundational frameworks, we do not cover these extensions in detail.
@@ 2.1 The Potential Outcomes Framework
The potential outcomes framework provides an intuitive and mathematically tractable foundation for causal inference. It was originally introduced by Neyman (1923) for randomized experiments and was later formalized and extended to observational studies by Rubin (1974). This was further extended by Robins (1986) to studies with time-varying treatments and confounders, as well as to direct and indirect effects and feedback of one cause on another. This framework is closely related to an independent line of work in econometrics that developed the so-called switching regression model (Roy, 1951; Manski, 1993). These developments allow researchers to pose well-defined causal questions, even in the absence of randomized experiments, by explicitly invoking assumptions that link observed data to potential outcomes. We refer interested readers to Rubin (2005); Morgan and Winship (2014); Imbens and Rubin (2015); Hernán and Robins (2025) for a more detailed introduction.
@@
The variables $Y(0)$ and $Y(1)$ are two distinct random variables representing hypothetical outcomes under different treatment assignments. They are not different realizations of the same variable. For each unit, only one of these two outcomes is realized and observed, depending on the actual treatment received. The other remains counterfactual. This leads to the Fundamental Problem of Causal Inference (Holland, 1986): we can never simultaneously observe both potential outcomes $Y(0)$ and $Y(1)$ for the same individual. From this perspective, the potential outcomes framework views causal inference as a missing data problem, since at least one of the two potential outcomes is missing for each unit.
@@
Potential outcomes are sometimes referred to as counterfactuals in the literature. Before treatment assignment, both $Y(1)$ and $Y(0)$ are potential outcomes. After treatment assignment, the outcome corresponding to the received treatment is factual, while the other is counterfactual. (However, many papers use the terms potential outcomes and counterfactuals interchangeably.) To avoid ambiguity, we use the term potential outcomes to collectively refer to both $Y(1)$ and $Y(0)$.
@@
The estimands we have introduced so far are sometimes referred to as interventional estimands, as they pertain to the distribution of outcomes under hypothetical interventions, such as setting the treatment to a fixed value. Another type of question that may be of interest involves counterfactual reasoning. For example, when $Y$ is binary, the persuasion rate is defined as $\mathbb{P}\{Y(1)=1\mid Y(0)=0\}$ (Jun and Lee, 2023), which quantifies the probability that an individual would take an action under treatment given that they would not have done so under control. Other examples include the probability of necessity, $\mathbb{P}\{Y(0)=0\mid A=1,Y=1\}$; the probability of sufficiency, $\mathbb{P}\{Y(1)=1\mid A=0,Y=0\}$; and the probability of necessity and sufficiency, $\mathbb{P}\{Y(1)=1,Y(0)=0\}$ (Pearl, 2009, § 9.2). Identification of counterfactual estimands is typically much more involved, as it relies on the dependence between potential outcomes, which cannot be recovered from observed data, even in randomized controlled experiments. We refer readers to Peters et al. (2017, Example 3.4) for an interesting example and discussions on whether counterfactuals or interventional quantities should serve as the basis for legal decisions; see also Robins and Greenland (1989), Robins and Greenland (1991), Pearl (2009, §1.4.4), and Wu and Wang (2026). Along with association questions, interventional and counterfactual questions are referred to as the three levels of the ladder of causation by Pearl and Mackenzie (2018, § 1).
A crucial component of the Potential Outcome framework is the Stable Unit Treatment Value Assumption (SUTVA) (Rubin, 1980), which posits that (i) there are no multiple versions of the treatment, and (ii) one unit’s outcome is unaffected by the treatment assigned to another unit, meaning that there is no interference between units.
The first part of SUTVA is closely related to the consistency assumption (Cole and Frangakis, 2009; VanderWeele, 2009; Pearl, 2010b), which is the fundamental assumption that links the observational world to the potential outcomes in a hypothetical interventional world. The consistency assumption states that if an individual receives treatment level $A=a$, then their observed outcome $Y$ equals their potential outcome under that treatment:
@@
A classic example of an ill-defined treatment arises when the so-called “treatment” is the causal effect of obesity (defined as BMI >30>30) on heart disease (Hernán and Robins, 2025).
@@
The consistency assumption discussed here is not the same as the notion of consistency used to describe a statistical property of an estimator (e.g., Lehmann and Casella, 1998), which refers to convergence in probability to the true parameter value.
The second part of SUTVA, namely the no interference assumption, implies that each unit’s potential outcome depends only on its own treatment assignment and not on the treatment received by others. This assumption can be violated in many real-world contexts. For example, in vaccine studies, the likelihood of an individual contracting an infectious disease may depend not only on whether they are vaccinated but also on the vaccination status of others in their community. Such interference is common in infectious disease epidemiology and social network settings, where spillover, peer effects, or herd immunity can induce dependencies across units. A growing body of work has extended the potential outcomes framework to accommodate interference. We refer interested readers, for example, to Kuang et al. (2020, §9) and Bajari et al. (2023) for surveys of this area.
@@
This condition implies that, conditional on $L$, the treatment assignment $A$ is as good as random and thus independent of the potential outcomes. Under this assumption and a positivity condition that ensures all levels of treatment are possible across values of $L$, the ACE is identified by the following formula:
@@
where the outer expectation is over the marginal distribution of $L$. Equation (2) was termed “direct standardization” in the early epidemiological literature (Neison, 1844) and is also a special case of the g-formula of Robins (1986). This expression can be estimated using regression, stratification, matching, inverse probability weighting, or doubly robust methods, depending on the structure and quality of the observed data.
@@ 2.2 Structural Equation Models
An NSPEM (Haavelmo, 1943; Strotz and Wold, 1960; Pearl, 2009; Spirtes et al., 2000; Halpern and Pearl, 2000) consists of a finite set of random variables $V=\{V_{1},\ldots,V_{p}\}$, a collection of exogenous (unobserved) random variables $\varepsilon=(\varepsilon_{1},\ldots,\varepsilon_{p})$, and a collection of measurable functions $f_{1},\ldots,f_{p}$, such that each endogenous variable $V_{j}$ is determined by an equation of the form
@@
Together, the functions $f_{j}$ and exogenous variables $\varepsilon_{j}$ specify a data-generating process and thereby induce a joint distribution over $V$. Importantly, NPSEMs do not need to be acyclic: cycles or feedback relations can in principle be allowed. When the system is acyclic, however, there exists (at least) one ordering $V_{\pi(1)},\ldots,V_{\pi(p)}$ such that each $V_{\pi(j)}$ depends only on variables earlier in the ordering. In this case, the structure of dependence among variables can be represented by a directed acyclic graph (DAG) (see Section 3.2). The model is then fully defined by the tuple $(f_{1},\ldots,f_{p},\varepsilon_{1},\ldots,\varepsilon_{p})$ together with a valid causal ordering (e.g., as obtained from Simon’s causal ordering algorithm (Simon, 1953)). The nonparametric nature of NPSEMs arises from the fact that no parametric form is assumed for the functions $f_{j}$, allowing for considerable flexibility in modeling complex causal relationships. A key assumption in many NPSEM applications is that the exogenous variables in $\varepsilon$ are jointly independent, yielding the special case known as the NPSEM with Independent Errors (NPSEM-IE).
@@
The Causally Interpretable Structured Tree Graph (CISTG) as detailed as the data (Robins, 1986) is mathematically equivalent to an NPSEM model (with no assumption on the errors); see Equations (10) and (14) below.
@@
The key assumption in the SEM framework that links the observational world to interventional worlds is the autonomy assumption (Aldrich, 1989), also known as the invariance or modularity assumption (Peters et al., 2017). It asserts that each structural equation represents an independent, self-contained causal mechanism that remains invariant under interventions on other variables. Formally, an intervention on a variable $A$ corresponds to removing the structural equation for $A$ and replacing it with a constant assignment. This is denoted by Pearl’s do-operator (Pearl, 2009), written as $\text{do}(A=a)$. Historically, Robins (1986) used the notation $g=a$ to express the same idea; Pearl’s do notation is now more widely used.
@@
Autonomy ensures that interventions can be represented by replacing the structural equation for the intervened variable, without having to re-specify the rest of the system. This feature allows us to predict counterfactual outcomes under hypothetical interventions. In real-world systems, however, autonomy is not always guaranteed (e.g., Cartwright, 2007). Administering the treatment may change the patient’s behavior or alter how the outcome is measured. For instance, it may modify the reporting mechanism or induce side effects that feed back into the outcome-generation process in an unmodeled way. In such cases, the function $f_{Y}$ may no longer be invariant to interventions on $A$. In other words, the mapping from $(L,A,\varepsilon_{Y})$ to $Y$ may differ depending on whether $A$ is set naturally or via intervention.
@@
The term Structural Causal Models (SCMs) is often used in the literature. It may refer to the NPSEM (e.g., Pearl, 2009, p. 203, Definition 7.1.1) alone or specifically to the NPSEM with independent errors (NPSEM-IE) (e.g., Pearl, 2009, p. 44, Definition 2.2.2); see also Peters et al. (2017, p. 62, Definition 6.2). Following Shpitser et al. (2022), we use our terminology here to emphasize the distinction between NPSEM and NPSEM-IE.
@@
The independent error and autonomy assumptions in the NPSEM-IE framework imply the principle of independent mechanisms (PIM) (Janzing and Schlölkopf, 2010; Peters et al., 2017). This principle states that the causal generative process can be decomposed into independent modules, one for each variable, that remain invariant to changes in the distributions of other variables. In the special case of two variables, PIM reduces to the so-called independence of cause and mechanism, which asserts that the conditional distribution $P(Y\mid X)$ is independent of the marginal distribution $P(X)$. In other words, under NPSEM-IE assumptions, the causal mechanism mapping $X\mapsto Y$ is invariant to changes in the distribution of $X$.